# MaxQ: Multi-Axis Query for N:M Sparsity Network

### 12 Dec 2023 v1

@@ Abstract
N:M sparsity has received increasing attention due to its remarkable performance and latency trade-off compared with structured and unstructured sparsity. However, existing N:M sparsity methods do not differentiate the relative importance of weights among blocks and leave important weights underappreciated. Besides, they directly apply N:M sparsity to the whole network, which will cause severe information loss. Thus, they are still sub-optimal. In this paper, we propose an efficient and effective Multi-Axis Query methodology, dubbed as MaxQ, to rectify these problems. During the training, MaxQ employs a dynamic approach to generate soft N:M masks, considering the weight importance across multiple axes. This method enhances the weights with more importance and ensures more effective updates. Meanwhile, a sparsity strategy that gradually increases the percentage of N:M weight blocks is applied, which allows the network to heal from the pruning-induced damage progressively. During the runtime, the N:M soft masks can be precomputed as constants and folded into weights without causing any distortion to the sparse pattern and incurring additional computational overhead. Comprehensive experiments demonstrate that MaxQ achieves consistent improvements across diverse CNN architectures in various computer vision tasks, including image classification, object detection and instance segmentation. For ResNet50 with 1:16 sparse pattern, MaxQ can achieve 74.6% top-1 accuracy on ImageNet and improve by over 2.8% over the state-of-the-art.
@@ 1 Introduction
Figure 1: Comparison of the accuracy-sparse pattern Pareto curve of the ResNet50 on ImageNet. MaxQ shows the top-performing Pareto frontier compared with previous N:M sparsity methods.
Among the many compression methods, network sparsity stands out as a highly effective tool for achieving practical memory and FLOPs saving. Most researchers have paid their attention to structured sparsity [18, 19, 25] and unstructured sparsity [14, 24]. However, both of them have certain shortcomings in terms of performance and acceleration, limiting their application. How to realize a better trade-off remains an open problem, as it hinges on factors like the granularity of sparsity and hardware design.
@@
Recently, N:M sparsity [29], a fine-grained sparsity pattern, is considered a highly promising solution. It can achieve a better trade-off between performance and latency compared to structured and unstructured sparsity thanks to its hardware and software co-design. Several studies have enhanced the performance of N:M sparsity. The pioneer work ASP [29] proposes to remove the two weights with the smallest magnitude from every four consecutive weights and use the pretrain-prune-finetune (PPF) pipeline to get the sparse network. This approach effectively preserves the performance of the dense network while it still suffers from expensive computation. To address this, Zhou et al. [1] proposes to learn N:M sparsity from scratch by extending a regularization term to Straight-Through Estimator, which improves the efficiency of sparse architecture update. Zhang et al. [37] treats the N:M sparsity problem as a combinatorial problem and solves it in an efficient divide-and-conquer manner. Although these methods improve the training efficiency or performance of N:M sparsity network, they are typically constructed upon the premise of weight block independence and do not differentiate the weight importance among blocks. At the same time, these methods apply N:M sparsity to the whole network N:M from the beginning of training, leading to critical information loss. As a result, they are still sub-optimal.
@@ 2.1 Sparsity Granularity in Network Compression
Mask Criteria. Extensive studies have investigated how to identify the critical weights in a network. Among them, the most widely used is the magnitude-based criterion that selects the pruning targets by their absolute value or $\ell_{1}$-norm. Apart from this, He et al. [19] proposed to prune filters via geometric median to compress CNN models with redundancy rather than those with ”relatively less” importance. Molchanov et al. [30] proposed to estimate the contribution of a neuron (filter) to the final loss via the first and second-order Taylor expansions. Lee et al. [24] introduced a saliency criterion based on connection sensitivity to identify structurally important connections in the network. Some other criteria, such as BN-based [28], have also proven effective.
@@ 3.1 Preliminaries
Here, we define the N:M sparsity problem. Without loss of generality, we suppose an L-layer CNN with parameters $\mathbf{W}=\{\mathbf{w}^{1},...,\mathbf{w}^{L}\}$, with $\mathbf{w}^{l}\in\mathbb{R}^{C^{l}_{\text{out}}\times C_{\text{in}}^{l}\times K^{l}_{\text{h}}\times K^{l}_{\text{w}}}$. $C^{l}_{\text{out}}$, $C^{l}_{\text{in}}$, $K^{l}_{\text{h}}$ and $K^{l}_{\text{w}}$ represent the number of output channels, input channels, kernel height and kernel width for $l$-th layer respectively. N:M sparsity groups every M consecutive weights along the input channel in each layer and requires each group to at most N non-zero elements. With this pattern, only N non-zero values in each group need to be stored and metadata is adopted to encode the position of each non-zero value. This process can compress the origin matrix and be accelerated by designated processing units (e.g. NVIDIA Ampere Tensor Cores). To this end, we first rearrange $\mathbf{w}^{l}$ to ${\mathbf{m}^{l}\in\mathbb{R}^{G^{l}\times M}}$, where $G^{l}=\frac{C^{l}_{\text{out}}\times K^{l}_{\text{h}}\times K^{l}_{\text{w}}\times C^{l}_{\text{in}}}{M}$ and denote $\mathbf{M}=\{\mathbf{m}^{1},...,\mathbf{m}^{L}\}$. Meanwhile, N:M binary masks $\mathbf{b}^{l}\in\left\{0,1\right\}^{G^{l}\times M}$ are introduced to achieve such sparsity Then, the optimization objective can be formulated as:
@@ 3.2 Multi-Axis Query
First, for $l$-th layer, we init all elements in $\textbf{b}^{l}$ to 1. For $t$-th epoch, we sort each element in $\mathbf{m}^{l}_{g}$ according to it absolute value and sort the $\mathbf{m}^{l}$ according to it’s $\ell_{1}$-norm to identify the set of blocks to apply N:M sparsity:
@@ 4.6 Performance Analysis
Figure 6: Parameter distribution from SR-STE at inference.
Figure 7: Parameter distribution from MaxQ at inference.
Training. The training efficiency of the algorithm is also an important aspect to consider. While many algorithms have smaller FLOPs theoretically, they may not be highly parallelized or have a large amount of memory access. Thus, there will be an increase but not a reduction in GPU days. To investigate the efficiency of various algorithms for N:M sparsity, we present the train speed, top-1 accuracy, and training FLOPs for SR-STE, LBC, and MaxQ in Tab. 9. For a fair comparison, their train speeds are tested with the same training script on the same machine with an NVIDIA RTX 3090 and automatic mixed precision. SR-STE is the fastest. MaxQ, although about 15% slower than SR-STE, achieves the highest top-1 accuracy among the three methods. The decrease in train speed primarily results from the multi-axis query, which incurs high memory access but involves minimal computational workload. Furthermore, although LBC has the theoretically least training FLOPs, it is the slowest among the three methods. Compared to the train speed of the dense network, LBC is almost 50% slower for a batch size of 128 and 30% slower for a batch size of 256. In summary, MaxQ stands out as the most efficient algorithm, striking the best balance between training speed and top-1 accuracy.
4.7 Quantization

### 17 Mar 2024 v2

@@ Abstract
N:M sparsity has received increasing attention due to its remarkable performance and latency trade-off compared with structured and unstructured sparsity. However, existing N:M sparsity methods do not differentiate the relative importance of weights among blocks and leave important weights underappreciated. Besides, they directly apply N:M sparsity to the whole network, which will cause severe information loss. Thus, they are still sub-optimal. In this paper, we propose an efficient and effective Multi-Axis Query methodology, dubbed as MaxQ, to rectify these problems. During the training, MaxQ employs a dynamic approach to generate soft N:M masks, considering the weight importance across multiple axes. This method enhances the weights with more importance and ensures more effective updates. Meanwhile, a sparsity strategy that gradually increases the percentage of N:M weight blocks is applied, which allows the network to heal from the pruning-induced damage progressively. During the runtime, the N:M soft masks can be precomputed as constants and folded into weights without causing any distortion to the sparse pattern and incurring additional computational overhead. Comprehensive experiments demonstrate that MaxQ achieves consistent improvements across diverse CNN architectures in various computer vision tasks, including image classification, object detection and instance segmentation. For ResNet50 with 1:16 sparse pattern, MaxQ can achieve 74.6% top-1 accuracy on ImageNet and improve by over 2.8% over the state-of-the-art. Codes and checkpoints are available at https://github.com/JingyangXiang/MaxQ.
@@ 1 Introduction
Figure 1: Comparison of the accuracy-sparse pattern Pareto curve of the ResNet50 on ImageNet. MaxQ shows the top-performing Pareto frontier compared with previous N:M sparsity methods [25, 1, 44, 45].
Among the many compression methods [22, 15, 24, 33, 37], network sparsity stands out as a highly effective tool for achieving practical memory and FLOPs saving. Most researchers have paid their attention to structured sparsity [20, 21, 30] and unstructured sparsity [16, 28]. However, both of them have certain shortcomings in terms of performance and acceleration, limiting their application. How to realize a better trade-off remains an open problem, as it hinges on factors like the granularity of sparsity and hardware design.
@@
Recently, N:M sparsity [35], a fine-grained sparsity pattern, is considered a highly promising solution. It can achieve a better trade-off between performance and latency compared to structured and unstructured sparsity thanks to its hardware and software co-design [35, 29]. Several studies have enhanced the performance of N:M sparsity. The pioneer work ASP [35] proposes to remove the two weights with the smallest magnitude from every four consecutive weights and use the pretrain-prune-finetune (PPF) pipeline to get the sparse network. This approach effectively preserves the performance of the dense network while it still suffers from expensive computation. To address this, Zhou et al. [1] proposes to learn N:M sparsity from scratch by extending a regularization term to Straight-Through Estimator [3], which improves the efficiency of sparse architecture update. Zhang et al. [44] treats the N:M sparsity problem as a combinatorial problem and solves it in an efficient divide-and-conquer manner. Although these methods improve the training efficiency or performance of N:M sparsity network, they are typically constructed upon the premise of weight block independence and do not differentiate the weight importance among blocks. At the same time, these methods apply N:M sparsity to the whole network N:M from the beginning of training, leading to critical information loss. As a result, they are still sub-optimal.
@@ 2.1 Sparsity Granularity in Network Compression
Mask Criteria. Extensive studies have investigated how to identify the critical weights in a network. Among them, the most widely used is the magnitude-based criterion that selects the pruning targets by their absolute value or $\ell_{1}$-norm. Apart from this, He et al. [21] proposed to prune filters via geometric median to compress CNN models with redundancy rather than those with "relatively less" importance. Molchanov et al. [36] proposed to estimate the contribution of a neuron (filter) to the final loss via the first and second-order Taylor expansions. Lee et al. [28] introduced a saliency criterion based on connection sensitivity to identify structurally important connections in the network. Some other criteria, such as BN-based [34], have also proven effective.
@@ 3.1 Preliminaries
Here, we define the N:M sparsity problem. Without loss of generality, we suppose an L-layer CNN with parameters $\mathbf{W}=\{\mathbf{w}^{1},...,\mathbf{w}^{L}\}$, with $\mathbf{w}^{l}\in\mathbb{R}^{C^{l}_{\text{out}}\times C_{\text{in}}^{l}\times K^{l}_{\text{h}}\times K^{l}_{\text{w}}}$. $C^{l}_{\text{out}}$, $C^{l}_{\text{in}}$, $K^{l}_{\text{h}}$ and $K^{l}_{\text{w}}$ represent the number of output channels, input channels, kernel height and kernel width for $l$-th layer respectively. N:M sparsity groups every M consecutive weights along the input channel in each layer and requires each group to have at most N non-zero elements. With this pattern, only N non-zero values in each group need to be stored and metadata is adopted to encode the position of each non-zero value. This process can compress the origin matrix and be accelerated by designated processing units (e.g. NVIDIA Ampere Tensor Cores). To this end, we first rearrange $\mathbf{w}^{l}$ to ${\mathbf{m}^{l}\in\mathbb{R}^{G^{l}\times M}}$, where $G^{l}=\frac{C^{l}_{\text{out}}\times K^{l}_{\text{h}}\times K^{l}_{\text{w}}\times C^{l}_{\text{in}}}{M}$ and denote $\mathbf{M}=\{\mathbf{m}^{1},...,\mathbf{m}^{L}\}$. Meanwhile, N:M binary masks $\mathbf{b}^{l}\in\left\{0,1\right\}^{G^{l}\times M}$ are introduced to achieve such sparsity Then, the optimization objective can be formulated as:
@@ 3.2 Multi-Axis Query
First, for $l$-th layer, we initialize all elements in $\textbf{b}^{l}$ to 1. For $t$-th epoch, we sort each element in $\mathbf{m}^{l}_{g}$ according to it absolute value and sort the $\mathbf{m}^{l}$ according to it’s $\ell_{1}$-norm to identify the set of blocks to apply N:M sparsity:
@@ 4.6 Performance Analysis
Figure 6: Parameter distribution from SR-STE at inference.
Figure 7: Parameter distribution from MaxQ at inference.
@@
Training. The training efficiency of the algorithm is also an important aspect to consider. While many algorithms have smaller FLOPs theoretically, they may not be highly parallelized or have a large amount of memory access. Thus, there will be an increase but not a reduction in GPU days. To investigate the efficiency of various algorithms for N:M sparsity, we present the train speed, top-1 accuracy, and training FLOPs for SR-STE, LBC, and MaxQ in Tab. 9. For a fair comparison, their train speeds are tested with the same training script on the same machine with an NVIDIA RTX 3090 and automatic mixed precision. SR-STE is the fastest. MaxQ, although about 15% slower than SR-STE, achieves the highest top-1 accuracy among the three methods. The decrease in train speed primarily results from the multi-axis query, which incurs high memory access but involves minimal computational workload. Furthermore, although LBC has the theoretically least training FLOPs, it is the slowest among the three methods. Compared to the train speed of dense network, LBC is almost 50% slower for a batch size of 128 and 30% slower for a batch size of 256. In summary, MaxQ stands out as the most efficient algorithm, striking the best balance between training speed and accuracy.
4.7 Quantization
@@ 5 Conclusion
Acknowledgements  This work is funded in part by the Key Research and Development Project of Zhejiang Province under Grant 2021C01035.